# Event-Based Simulation of Stochastic Memristive Devices for Neuromorphic Computing

### v1

@@ II-D Switching Mechanism
    $E_{a}=qV_{a}$ is the activation energy (the energy barrier height). This increases or decreases both the up and down switching rates.
@@
    $E_{\text{off}}=qV_{\text{off}}$ (added, rather than subtracted, for $\lambda^{\prime}_{down}$) is the energy offset that leads to a shift of the equilibrium point from equal numbers of states being in the high or low resistive state. This adjusts the balance of switching rates under zero bias.
@@
    $k$ is a constant factor used to convert an applied bias to the effective voltage used for determination of the effect voltage and subsequent shift in the barrier height, accounting for material-specific response factors, such as lattice periodicity.
@@
    $V(t)$ is the voltage across the device at time $t$, which will be added for $\lambda^{\prime}_{down}$ as opposed to being subtracted. This gives an additional energy offset when scaled by $q$.
  5. 5.
    $T(t)$ is the time-dependent temperature variable in degrees Kelvin. This may be influenced by the history of applied voltages and the device state history, which directly impact the rate of switching, as well as the equilibrium point when $E_{\text{off}}$ is not 0.
  6. 6.
    $\rho(t)$ has a similar impact to the temperature and is the volatility state variable, that may also be dependent on the time and voltage/state history.
In this model, application of a voltage bias, $V(t)$, lowers or raises the activation energy barrier in one direction, simultaneously raising or lowering it in the other direction. This is illustrated in Figure 3. Note also that, for the purposes of simplicity, in this paper, we assume, without loss of generality, that $k=1$.
Figure 3: Illustration of the effect of the different parameters on the Boltzmann barrier heights in the forward and reverse switching directions. In a) we see the standard barrier, unmodulated by an offset or an applied voltage. b) shows the barrier modified by an energy offset $E_{\text{off}}$. c) shows the energy barrier modulated by the offset in addition to the modulation of the barrier resulting from the application of a voltage bias, $E_{v}$. The relative energy barrier height for each local minimum determines the respective Poisson switching rates $\lambda^{\prime}_{\text{up}}$ and $\lambda^{\prime}_{\text{down}}$.
The volatility state variable is an additional state variable that is introduced to incorporate stateful effects on the evolution of the resistive state, discussed further in Section II-E.
II-D1 Boltzmann Activation Function
To determine the rates of the Poisson processes, we made use of a Boltzmann occupation function, rather than the Fermi-Dirac occupation function (logistic function) used to determine the switching probabilities in the GMSM. We make use of the Boltzmann distribution due to it being a good model of probabilistic behaviour in microscopic thermodynamic systems, modulated by energy barriers.
In the case of a metal oxide memristor, we envision ions hopping over energy barriers, mediated by voltage bias and temperature conditions, resulting in the movement of ions into more or less favourable locations for conduction (modulating the instantaneous resistance). The nature of the ions is not crucial: for example, they could be metal ions (as in ECM devices), or oxygen vacancies (as in VCM devices). There are many possible mechanisms that could drive the movement of ions in an ionic movement memristive device. Filamentary, interfacial, and bulk switching (e.g. Mott insulator-metal transition through ionic doping/trapping) mechanisms have been proposed. Such a physical switching mechanism motivates the use of a Boltzmann modelling approach. The use of sigmoid-like or Boltzmann-like activations is common in models of memristor state evolution, and a sigmoid activation is used in the GMSM simply to bound the probability of state transitions between 0 and 1 [1].
$\sinh$ switching behaviour (for an applied voltage) naturally arises out of modelling that uses an energy-barrier assumption with a height and hopping rate that is exponential in the applied voltage. $\sinh$-switching behaviour is something that has been observed in a variety of memristive devices and attributed to phenomena such as exponential ionic drift for high fields, or Joule heating effects [12, 13]. This can be observed by subtracting the instantaneous rate of upward transitions (transitions from the down state to the up state) from the downward transitions (transitions from the up state to the down state) as follows.
In [13], Strukov & Williams note that for the rigid point ion model, the energy barriers in the forward and reverse directions for the drift of ions in a lattice would be distorted, through the application of an electric field, by ±$\pm q\mathcal{E}a/2$, where $\mathcal{E}$ is the field, and $a$ is the periodicity of the lattice. The characteristic length $a$ need not be lattice periodicity, therefore we simply combine the impact of the lattice periodicity, specific to the rigid point ion model, and the electric field into a parameter $V=\mathcal{E}a$, thus the distortion will be ±$\pm qV/2$, respectively.
The average energy of any ion will be approximately the thermal voltage $V_{T}=k_{B}T$, thus providing a mechanism for thermally accelerated drift, through Joule heating effects.
Based on this, if we assume that the rate of hopping upwards or downwards for each “switch” is given by a rate equation based on the Boltzmann probability density [^1] The Boltzmann probability density is a common model for the probability density of particles at thermal equilibrium., then we can say that the rate of transition from the high resistance state to the low resistance state for individual switches is given by:
$r_{\text{up}}\propto e^{-(E_{a}-qV/2)/k_{B}T}$
while for the transition from the low resistance state to the high resistance state we have:
$r_{\text{down}}\propto e^{-(E_{a}+qV/2)/k_{B}T}$
In turn, we can see that subtracting the rates gives us the instantaneous, net rate for movement from a high resistance to a low resistance state:
$r_{\text{net}}$ $\propto e^{-(E_{a}-qV/2)/k_{B}T}-e^{-(E_{a}+qV/2)/k_{B}T}$ (5)
$\propto e^{-E_{a}/k_{B}T}\sinh{(-qV/2k_{B}T)}$ (6)
This is similar to the equation for high field ionic drift for the rigid point ion model described in [13]. However, the net rate is only an instantaneous rate of change. In our model, any event (transition from up to down, or down to up) causes a change in the number of switches in each state, and thus a linear and proportional change to the rates $\lambda_{\text{up}}$ and $\lambda_{\text{down}}$. As a result, we must re-calculate the rates whenever there is a change in the state.
II-E Volatility State Variable
Memristive systems, such as PCM and Metal Oxide devices, differ significantly in their structure, material composition, and switching/conduction mechanics, but can nevertheless exhibit similar characteristics. For example, in the case of their switching behaviour under bias, both types of device have associated physical phenomena which cause the device’s state to autonomously change under zero bias, following initial programming. This motivates the introduction of a unified way of describing these disparate physical phenomena for the purposes of modelling.
We introduce the notion of “volatility”. The term “volatility” is sometimes used to refer to the phenomenon of resistive drift - the tendency for the resistance of a device to deviate from its initial value over time. We here generalise the term to include any kind of transient physical phenomenon that can affect the switching or conduction of a memristive device. Examples of physical phenomena that fall under the umbrella of volatility include:
  1. 1.
    Joule heating resulting in nonlinear switching behaviour and/or accelerated carrier drift under bias in various memristive device types, including metal oxide and PCM devices.
  2. 2.
    Structural relaxation in PCM devices.
  3. 3.
    Structural disruption that causes reconfiguration in diffusive memristive devices.
The presence of volatile mechanics in a device tends to express itself in several ways: memory-dependence in switching and conduction behaviour; self-excitatory or self-inhibitory behaviour during switching and conduction leading to non-linear switching and/or conduction; and autonomous state evolution as a result of temporary changes within the devices, such as increased temperature or disorder in the material. The unification of the phenomena removes the need to identify the exact cause of memory-dependent switching behaviour in memristive devices, allowing us to focus only on the consequences for the evolution of the state variable. The volatility state variable summarises the device history that is relevant to future switching behaviour into a single state variable.
In the general sense of the word, an increase in volatility is associated with an increase in the state transition probabilities (in the discrete setting). This contrasts with a directed bias, which will result in an increase in one or more switching rates at the expense of decreasing one or more other switching rates.
Intuitively, if we imagine switching events to be associated with a change in the state of the device, an increase (decrease) in volatility means that the number of state change events per unit time increases (decreases). Depending on the degree to which the different switching rates increase, relative to one another, a change in the volatility may also have an impact on the local metastable (equilibrium or capturing) state(s).
Note that our definition of volatility motivates the way in which we model its effect mathematically for a given switching probability model. For example, in later sections we will explore the use of a Boltzmann activation energy model, and the volatility will have a similar effect to the temperature, increasing all switching rates simultaneously.
II-E1 Phenomena Captured by the Volatility State Variable
One key observation is that volatility is higher immediately after a programming event in a range of memristor devices, including ECM and VCM ionic movement memristors, and in PCM devices. It may also be the case that the degree of volatility is dependent in some way on the nature of the programming (i.e., it is not a memoryless phenomenon given the memristive state alone), with more extensive programming pulse application or a higher deviation from a previous stable level resulting in a higher degree of volatility. Volatility then tends to relax to a stable level in the absence of any stimulating voltage input (or change in ambient conditions such as temperature).
In the examples below, we detail experimental evidence that further justifies and supports the theoretical motivation for using a framework involving volatility as a state variable.
Joule Heating
In the case of filamentary switching behaviour in metal oxide memristors, the drift of oxygen vacancies is often non-linear (exponential) with respect to the applied bias. This is attributed to multiple possible phenomenon, including high-magnitude local electric fields, and possible accelearation through Joule heating effects within the device. It is therefore the case that Joule heating may produce memory-dependent switching behaviour, with the internal temperature of the device - itself affected by the history of voltage applied across the device - affecting the rate of switching.
When the device is being programmed, energy is dissipated and held as heat energy in the device. Over time, if there is no stimulating voltage, Newton’s law of cooling applies and the device returns to the ambient temperature. This is a dynamic process, and the cooling may not be fully complete before subsequent programming pulses. Cooling also occurs during the application of programming pulses, with cooling happening more rapidly, the higher the internal temperature of the device.
The actual picture governing volatility may not be fully described through the consideration of Joule heating effects. Wang et al. have examined the properties of temperature and diffusive dynamics for diffusion-driven, volatile tantalum-oxide memristors [14]. They state that temperature cannot be the driving factor in state relaxation for these diffusive, uni-polar devices, since temperature increases produced by Joule heating relax on the order of $1/\kappa$, whereas the memristor relaxation (stochasticity) occurs on a much longer order of $10-201/\kappa$. Thus, they hypothesise, diffusion is the main driving factor, rather than a remnant increase in temperature. This is applicable to the devices studied, which are based on the formation of conductive bridges formed from silver metal nanoparticles through $SiO_{x}N_{y}$ dielectric substrate doped with silver ions, and sandwiched between $Ag$ electrodes. Their conclusion is that Joule heating’s role in the switching process, or subsequent relaxation (switching in the absence of bias) can vary dependent on the inherent volatility of the device.
Structural Relaxation in PCM
Structural relaxation is a phenomenon in PCM (e.g. GST), that results from the gradual consumption of point defects in the amorphous structure over time, following heating followed by rapid cooling that results in a glass-like structure.
The number of unconsumed defects in the device can be thought of as the level of volatility, which is highest after programming when the device has not had time to partially relax back into a crystalline configuration. As the defects are consumed, the volatility decreases. A more detailed description is as follows: after programming, the device relaxes towards the ideal glass configuration, through the consumption of defects in amorphous GeTe that cause the amorphous phase to differ significantly from the crystalline phase. Although drift does not cause the phase to become crystalline, it is generally agreed that defects in the amorphous GeTe are consumed, moving the bond network towards ’structures with chemical order and coordination numbers similar to those of the crystalline phase’. Note that we emphasise that the structure is similar to the crystalline phase, but it is not identical. True relaxation to the crystalline phase would cause full restoration of the high resistive state [15], which is possible if the device is not programmed fully to the amorphous phase using partial SET operations [16].
Electrochemical Metallization: Entropy and Disruption of Structure
In electrochemical metallization (ECM) devices, the entropy of the oxygen vacancies in the device may also prompt a memory-dependent response. Minimisation of interfacial energy may prompt reconfiguration of metal ions within the insulator bulk for diffusive memristors. This phenomenon is described for the case of diffusive memristors in [14].
The general form of volatility is a state variable which directly modifies the global rate of switching, whose increase is a function of external inputs such as voltage, or current, and which has a characteristic decay time constant, $\tau_{x}$. Example implementations of volatility are given in Section IV-D3, where the effects of both Joule heating/Newton cooling, and of arbitrary volatile effects are considered.
II-F Practical Simulations
@@
II-F1 Point Process Sampling
@@
II-F2 Changes to the Distribution at Events
@@
We use the memoryless property of the exponential distribution, which states that if an event has not occurred by time $\tau$ then the distribution for an event to occur at time $t+\tau$ is still exponentially distributed. The only distinction is that we now modify the rate parameter for each distribution, which can be calculated as a simple linear scaling of the rate parameter, since the sum of N independent and identically distributed exponential distributions with rate parameter $\lambda$ is an exponential distribution with rate parameter $N\lambda$, as described in Section II-D.
II-F3 Simulation Algorithm
@@
In order to incorporate non-event based inputs, including variables that may be continuous functions of time such as temperature or applied bias, or piecewise constant functions, we can sample these using threshold-based, or event-driven, sampling. For example, the volatility state variable, introduced in Section II-E, is a continuous function of time. Given an independent input signal $m_{i}(t)$, we can define the associated marked point process, defined by the set of times $\mathcal{T}_{i}$, at which the events occur, to be:
@@
III Model Analysis
III-A Model Analysis
Our model has thus far been defined, but we would also like a computationally efficient way of evaluating it. Our model is a marked [^2] A point process whose events are associated with additional information, such as amplitude. In this case - the direction of switching of the state variable. Hawkes [^3] A point process whose events impact the rate parameter (traditionally self-excitatory, but also allowing for self-inhibitory behaviour). process with both self-excitatory and self-inhibitory behaviour. We would like to demonstrate different ways of looking at the behaviour of the model, in order to motivate the development of efficient schemes for numerical simulations and prediction of the statistics.
The majority of our analysis in this section will be under the assumption of static environmental conditions (temperature and bias), since under varying conditions, additional modifications to the switching parameters, and thus, to the equilibrium point of the model, will occur. The change of the memristive state in the absence of a voltage bias is usually referred to as the phenomenon of resistive drift, or volatility. The equilibrium point of the model under steady state conditions is identified in Section III-A1.
III-A1 Calculation of the Equilibrium Point (Metastable Point)

### v2

@@ II-C Volatility State Variable
    Joule heating in metal oxide and Phase Change Memory (PCM) devices.
@@
    Structural relaxation in PCM devices [11, 12].
@@
    Interfacial energy minimisation in diffusive memristive devices [13].
These phenomena are unified by their modulation of the switching characteristics under particular bias and temperature conditions. All of these phenomena can cause a temporary increase in the rate of change of the memristor state variable for a given voltage bias. This can manifest in the form of several switching phenomena:
  1. 1.
    Metastability - the tendency for a device to revert to a more stable neighbouring state after switching.
  2. 2.
    Resistive drift (autonomous state evolution) - a phenomenon involving a change of state under zero bias.
  3. 3.
    Non-Linear switching e.g. exponential ionic drift [14].
@@
    Accelerated switching behaviour, where the frequency or duration (spacing in time) of applied voltage pulses impacts the rate of switching (see Section IV-C1).
As such, we propose a unified modelling approach in the form of a “volatility state variable”. The term “volatility” is sometimes used, in the sense of instability, as a synonym for the phenomenon of resistive drift - the tendency for the resistance of a device to deviate from its initial value over time. We here use the term to instead refer to any kind of transient physical phenomenon that temporarily affects the rate of switching or conduction of a memristive device. This general description is of value from the perspective of modelling for neuromorphic computing applications, where we would like to understand the aforementioned behaviours such that we can control them to perform computation.
We define a volatility state variable, $\rho(t)$, as follows: one or more state variables which directly modify the global rate of switching (both $\lambda_{\text{up}}$ and $\lambda_{\text{down}}$), which has a characteristic time constant of decay, $\tau_{x}$ and will tend to decay to an equilibrium value $\rho_{\text{eq}}$, which may be modulated by external (exogenous) inputs of voltage or temperature:
$\frac{d\rho(t)}{dt}=\left(\rho_{\text{eq}}(v(t),T(t))-\rho(t)\right)/\tau_{\text{volatile}}$ (3)
The exact nature of the volatility model will be dependent on the type of device. Its role is also visualised in Figure 1.
II-D Practical Simulations
@@
II-D1 Point Process Sampling
@@
II-D2 Changes to the Distribution at Events
@@
We use the memoryless property of the exponential distribution, which states that if an event has not occurred by time $\tau$ then the distribution for an event to occur at time $t+\tau$ is still exponentially distributed. The only distinction is that we now modify the rate parameter for each distribution, which can be calculated as a simple linear scaling of the rate parameter, since the sum of N independent and identically distributed exponential distributions with rate parameter $\lambda$ is an exponential distribution with rate parameter $N\lambda$, as described in Section II-B.
II-D3 Simulation Algorithm
@@
In order to incorporate non-event based inputs, including variables that may be continuous functions of time such as temperature or applied bias, or piecewise constant functions, we can sample these using threshold-based, or event-driven, sampling. For example, the volatility state variable, introduced in Section II-C, is a continuous function of time. Given an independent input signal $m_{i}(t)$, we can define the associated marked point process, defined by the set of times $\mathcal{T}_{i}$, at which the events occur, to be:
@@
II-E Discussion
II-E1 Impact of the Parameter $N$
One strength of a discrete modelling strategy, with a finite number of possible system states (in our case, $N$), is that simulations can be conducted at different resolutions, depending on precision requirements, and the scale of interest. For example, if a memristive device is to be used as a binary storage device, where only the states $1$ and 00 are distinct, then a value of $N=2$ suffices to model the readout value of interest. In contrast, for a simulation where multiple stable states are to be for information storage, a larger value of $N$ may be required. The “best” value of $N$ to use is a matter of perspective - it depends on the extent to which regions of device resistance or behaviour are considered to be equivalent for the intended purpose, as exemplified by binary storage, where the many possible configurations of the device are summarised by a single bit (2 possible values) of information.
As a consequence of the fact that the variance and mean of a Poisson process are equal to the rate of the process, the magnitude of the parameter $N$ will scale the expected value and the variance of the final state in a linear fashion. As a result, the standard deviation in the final state, after a given delay, is proportional to $\sqrt{N}$. We can therefore modulate stochasticity by by changing $N$. When we increase N, the average jump time decreases linearly with N (the rates of up and down jumps increase proportional to N). This means that over the timescale of interest, many more events will tend to occur, with each trial also contributing to a smaller percentage change in the resistance, resulting in a decrease in the variance with increasing $N$. Continuous, deterministic switching behaviour can be approximated in the limit as $N\rightarrow\infty$.
II-E2 Elimination of Discrete-Time Sampling Error
We note that the approximation in Equation 1 is valid only for small values of $\Delta t$. If we let the value of $\Delta t$ grow large, then errors associated with discretising the timestep begin to emerge. The larger the value of $\Delta t$, the more significant a change will have occurred in the parameters used to calculate the transition probabilities, leading to modelling inaccuracies. This is a problem for any discrete-time approximation of a continuous-time problem i.e. any finite difference method for approximating a continuous-time differential equation. Error-free modelling can only be achieved in the limit as the timestep vanishes, i.e., $\Delta t\rightarrow 0$. The switching probability at each point in time will change with each switching event; however, updates are not instantaneously incorporated into the model, but are only applied every $\Delta t$ seconds, over which time many switching events may have occurred, resulting in an error. Such errors compound over time.
An advantage of our event-based adaptation of the GMSM is to eliminate the error associated with discrete time-stepped simulation. However, it should be noted that using a value of N that is too coarse will still introduce a quantisation error. This error does not, however, compound over time and the expected value of the process after given number of switching events will not be biased, since the value of $n_{eq}$ scales linearly with $N$, and the associated parameters of the readout equation can be adjusted accordingly.
Note that integrate and fire sampling for continuous functions of time, as described in Section II-D3, will unavoidably introduce discretisation errors. If the functions vary slowly in time, then the number of additional events due to the introduction of additional functions will be small. However, if the functions vary quickly in time and the number of generated events is large, then the benefits of event-based modelling approaches diminish, and the model’s events become dense in time. This still, however, offers benefits over a uniform sampling approach, since it associates the density of samples in time with the rate of information change that impact the model updates (and hence the error associated with discrete-time sampling).
III Metal Oxide Memristor Modelling
In this section, we look at the application of the modelling framework to the modelling of resistive drift phenomenon under zero bias in titanium dioxide memristors. As such, we give an illustrative example of the proposed event-based modelling framework, detailing a readout equation, switching probabilities, and the volatility state variable for this setting.
III-A Physical Model
Modelling of metal oxide Metal-Insulator-Metal (MIM), and more specifically - Titanium Dioxide MIM - has been previously modelled as a single conductive pathway of varying resistance [9], and other similar devices, such as tantalum oxide memristors have been modelled as exhibiting the dominance of a single filament or conductive region [15].
Based on initial investigation of a drift dataset (see Section III-E below), we model the conductance of the device as linearly proportional to state variable $n$. We hypothesise that the mechanism underlying the switching of the device modulates the width of a conductive region within the active substrate, bridging a connection between the electrodes. Phenomena that could account for this behaviour include a single conductive filament that grows or reduces in width, or multiple conductive filaments that each contribute linearly to the conductance of the device. We observe that it is necessary to use a value of $n_{\text{extra}}$ and $m_{\text{extra}}$ much greater than the value of $N$. This suggests a thresholding effect, where the rate of switching is relatively constant in the switching region, under a constant voltage bias. This corresponds somewhat to the ideal behaviour for a memristor, which is why such titanium dioxide devices were presented as candidate memristive devices originally [9]. We assume a large $n_{\text{extra}}=10\times N$ such that the switching rates remain relatively constant across the considered range of resistance values.
III-B Readout Equation
We can also invert the equation to find $R$ in terms of $n$:
III-C Switching Probability
To determine the rates of the Poisson processes, we make use of a Boltzmann occupation function that is dependent on the voltage bias and the temperature, as well as static parameters, rather than the Fermi-Dirac occupation function (logistic function) used to determine the switching probabilities in the GMSM. This is due to it being a good model of probabilistic behaviour in microscopic thermodynamic systems, modulated by energy barriers.
$\lambda^{\prime}_{\text{up}}(t)$ $=\exp{\left(-\frac{E_{a}-V(t)q/2-E_{\text{off}}/2}{k_{B}T(t)(1+\rho(t))}\right)},$ (6)
$\lambda_{\text{up}}(t)$ $=n(t)\lambda^{\prime}_{\text{up}}(t),$ (7)
with the parameters defined in Table I:
$E_{a}=qV_{a}$ Activation energy (the energy barrier height). This increases or decreases both the up and down switching rates.
$E_{\text{off}}=qV_{\text{off}}$ Energy offset that leads to a shift of the equilibrium point from equal numbers of states being in the high or low resistive state. This adjusts the balance of switching rates under zero bias. (Added, rather than subtracted, for $\lambda^{\prime}_{down}$).
$V(t)$ Voltage across the device at time $t$, which will be added for $\lambda^{\prime}_{down}$ as opposed to being subtracted. This gives an additional energy offset when scaled by $q$.
$T(t)$ Time-dependent temperature variable in degrees Kelvin. This may be influenced by the history of applied voltages and the device state history, which directly impact the rate of switching, as well as the equilibrium point when $E_{\text{off}}$ is not 0.
$\rho(t)$ Volatility state variable. Has a similar impact to the temperature in affecting the global switching rate (in fact, the temperature can be considered a volatility state variable as well).
TABLE I: Parameters and inputs to the Poisson rate parameter model for titanium dioxide memristive devices.
In this model, application of a voltage bias, $V(t)$, lowers or raises the activation energy barrier in one direction, simultaneously raising or lowering it in the other direction.
III-D Volatility
In the case of thin film memristors, volatility can be decomposed into two variables. The first of these is the temperature (that may increase due to Joule heating) and the second is a dimensionless volatility state variable to account for additional volatile effects, such as disruptions to the internal structure of the memristor on application of a bias voltage and subsequent restoration of order following its removal (due to interfacial energy minimisation).
III-E Dataset
To evaluate the behaviour of the model under controlled conditions, we created a dataset based on the evolution of memristive states [17]. We collected 2760 device hours of data across two separate memristor arrays. We set each of the devices in the array to a random initial resistance and measured the evolution of the resistance over time, through Kelvin sensing that used a 30ms pulse of 0.3V magnitude to measure the instantaneous resistance of the device, every 10 seconds, over varying numbers of hours. The devices used were thin-film Titanium Dioxide ($TiO_{2}$) memristors with $Al_{2}O_{3}$ interstitials (the same as [9]), placed on the diagonal of crossbar arrays to mitigate sneak path currents. The measurements were obtained using the ArCONE memristor characterisation platform [18].
Originally, a total of 219 data series were collected. After collection, the series were inspected and a data cleaning step was performed. As part of this step, series for memristors which had failed or degenerated to a non-volatile state after programming, as well as clearly anomalous measurements within series due to measurement errors, were removed. Following data cleaning, 23 series were excluded from the dataset, leaving 196 series.
III-E1 Quantisation
$\mathcal{R}_{\text{bound}}=$$\{-\infty,R(n_{\text{thresh}}+0.5),$
$R(n_{\text{thresh}}+1.5),\cdots,R(N-0.5),\infty\}$
$\mathcal{R}_{\text{val}}=$$\{R(n_{\text{thresh}}),R(n_{\text{thresh}}+1),$
$\cdots,R(N-1),R(N)\}$
III-E2 Pairs
III-F Model Fitting
III-F1 Calculation of the Equilibrium Point (Metastable Point)